# From Demonstrations to Safe Deployment: Path-Consistent Safety Filtering for Diffusion Policies

### v1

@@ VI Evaluation
    With our path-consistent safeguarding method, task success rate remains close to the unsafe policy ($\pm$ $5\text{\,}\%$) in our conducted experiments.
@@ VI-A Simulation Experiments
TABLE I: Robomimic task success rate results on 100100 rollouts. PACS achieves higher success rates than other safety filters.
@@
Contr. barrier func. [singletary2022safety] 0.11 0.00 0.00 0.04
@@
First, we evaluate the performance of PACS against an unsafe operational space controller and a control barrier function baseline [singletary2022safety] on the three robomimic [mandlekar2021matters] benchmark tasks Lift, Can, and Square. We integrate these tasks in human-robot-gym [thumm_2024_human_robot_gym] and add a spherical dynamic object that moves in pre-defined patterns on the surface of the table. The goal is to fulfill the original task while ensuring the safety constraints $c_{\text{safe},\text{SSM}}$ (for operational space controller, control barrier function, and SSM) or $c_{\text{safe},\text{PFL}}$ (for PFL). To evaluate H3, we further test a single-action version of PACS that sequentially computes new intended trajectories from each action in the action chunk.
We train the DP with a temporal UNet architecture [chi2023diffusion] on the provided dataset of $200$ proficient human teleoperated demonstrations per task. The policy receives the end effector and object poses of the last two timesteps as observations and predicts desired changes in the end effector pose and gripper position as actions. The operational space controller directly controls the robot in Cartesian space to follow these goals. For all other approaches, we use an inverse kinematics solver [coumans_2016_PyBulletPython] to translate the desired changes in end-effector pose into desired joint state actions and control the robot in joint space. Additional implementation details can be found in Table III. For the operational space controller baseline, we use $h=8$ action execution steps and apply each action for $\Delta t=$50\text{\,}\mathrm{ms}$\$. To evaluate the impact of the safety mechanism and the inverse kinematics translation separately, we also test the performance of our safety shield in OFF mode. Here, we compute the same intended trajectory in joint space, but never execute a failsafe trajectory.
@@
TABLE II: Constraint formulations for our real-world tasks.
@@ VI-B Real-World Setup
Second, we conduct hardware experiments with a Franka FR3 manipulator and consider the example application of HRI, as safety is particularly critical in this area.
@@ VI-B1 Tasks
We consider three tasks that require the robot to move safely in close proximity to humans, as visualized in Figure 3:
  • Sorting: We place two red and three green blocks randomly on the table. The robot must put the red blocks into a box while the human is taking the green blocks. To ensure safety, no collisions between the robot and the human are allowed ($c_{\text{safe},\text{SSM}}$).
@@ VI-B2 Safety Constraints
We use the SaRA framework [sven_2022_SaRA] for computing the reachable occupancies and intersection checking in real time. We formally identify the safe energy thresholds in (3) for the different tasks [thumm_2025_GeneralSafety]. Further details and numerical values are provided in Table II. We leverage Ruckig [Ruckig] to compute the intended trajectory (6), using the standard joint limits of the FR3, velocity limits between $1.25\text{\,}$ and $2\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-1}$, an acceleration limit of $10\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-2}$ and a jerk limit of $400\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-3}$ for each joint.
@@ VI-B3 Policies
We deploy two generative policies: DP [chi2023diffusion] that uses denoising diffusion [ho2020denoising] and SmolVLA [shukor2025smolvla] using flow matching [lipman2022flow]. The observations consist of joint angles and the images of a wrist and a third-person camera, and the actions are desired changes in joint angles and the gripper width. We provide details on the training data and the policy implementation in Table III. We use CRISP [pro2025crispcompliantros2] for data recording and for translating the joint commands to motor torques. The SmolVLA policy struggled with Sorting and Feeding, likely due to the high precision required; therefore, we only evaluated it on the Handover task.
@@ VI-C Real-World Results
We deploy the policies with our safety filter OFF and active, similar to the simulation experiments. As shown by the results in Table IV, the policies achieve very similar task success rates of about $80\text{\,}\%$ on average with and without our safety filter, supporting H2. The safeguarded policies do not exhibit constraint violations in our experiments. In contrast, the nominal policies are in an unsafe state $56\text{\,}\%$ of the time with each rollout containing constraint violations, resulting in a Safe Success of $0\text{\,}\%$, which confirms H1. These results highlight the necessity of online safety mechanisms when deploying DPs in dynamic environments and the effectiveness of PACS for this purpose.
TABLE IV: Impact of safeguarding in our real-world experiments. PACS guarantees safe deployment and maintains high task success, resulting in a high Safe Success rate.
@@
We also compare our path-consistent safety approach with a popular reactive method; control barrier functions [singletary2022safety]. Since control barrier functions are better suited for coexistence than for collaboration tasks, we conduct this comparison for the Sorting task. As can be seen from Table V, PACS achieves a $37\text{\,}\%$ higher success rate than the control barrier function-based filter. To further examine this inferior performance, we visualize the training data distribution and the end effector positions of randomly selected rollouts in Figure 4. While our method keeps the policy in distribution, the control barrier function often pushes the robot into out-of-distribution states from which the policy struggles to recover, leading to a higher failure rate. Considering our simulation results, we find that PACS overall achieves a $51\text{\,}\%$ higher task success rate than the control barrier function safety filter. Besides, we observe that even if the policy manages to recover, this usually takes a few seconds, leading to a higher task completion time on average. The average computation time of one safety step with PACS is $0.20\text{\,}\mathrm{ms}$ compared to $0.64\text{\,}\mathrm{ms}$ for the control barrier function, and recalculating the intended trajectory takes about $5\text{\,}\mathrm{ms}$. This supports H4 and shows that our method can be deployed in real time in highly dynamic environments.
@@
Figure 4: End effector paths for the Sorting task. The color intensity of the trajectories indicates the velocity, and the intensity of the shaded grey areas visualizes the training distribution. Our safety filter slows down the policy without leaving the desired path when the human is nearby. In contrast, the control barrier function (CBF) pushes the robot away from unsafe states, which often leads to out-of-distribution (OOD) states from which the policy cannot recover.
@@
TABLE V: Baseline comparison and speed analysis for the Sorting task. PACS achieves higher task performance than control barrier functions and can even perform the task faster than the nominal policy.
@@
Lastly, we test our safety filter for the Feeding task with a real human. As shown in our accompanying video, the robot slows down when approaching the human, and the safeguarded policy inserts and withdraws the fork carefully and precisely. These experiments highlight the potential of our approach to harness the expressiveness of DPs for highly safety-critical applications, such as healthcare.
@@ VII Conclusion
We propose PACS, a general framework for safely deploying action chunking-based imitation learning policies, such as DPs and vision-language-action models, in dynamic environments. Our insight is that safety interventions should be path-consistent to avoid out-of-distribution situations and maintain high task performance. We achieve this by computing an intended trajectory from the action chunk and monitoring it using reachability analysis. Our experiments show that PACS can guarantee safety in real time and achieve high task success even on challenging real-world HRI tasks, whereas reactive safety filtering with control barrier functions frequently leads the policy into unrecoverable states. Since we do not perform constraint-aware online replanning, which would require conditioning the policy on a constraint parameterization, our method is specifically designed to satisfy dynamic constraints, such as moving obstacles. We consider this work an important step towards deploying generative policies in safety-critical, human-centered environments.

### v2

@@ VI Evaluation
    With our path-consistent safeguarding method, the task success rate remains close to the unsafe policy ($\pm$ $5\text{\,}\%$) in our conducted experiments.
@@ VI-A Simulation Experiments
TABLE I: Robomimic task success rate results on 100100 rollouts.
@@
Contr. barrier func. [41] 0.11 0.00 0.00 0.04
@@
First, we evaluate the performance of PACS against an unsafe operational space controller and a control barrier function baseline [41] on the three robomimic [28] benchmark tasks Lift, Can, and Square. We integrated these tasks in human-robot-gym [45] and added a spherical dynamic object that moves in pre-defined patterns on the surface of the table. The goal is to fulfill the original task while ensuring the safety constraints $c_{\text{safe},\text{SSM}}$ (for operational space controller, control barrier function, and SSM) or $c_{\text{safe},\text{PFL}}$ (for PFL). To evaluate H3, we further tested a single-action version of PACS that sequentially computes new intended trajectories from each action in the action chunk.
We trained the DP with a temporal UNet architecture [12] on the provided dataset of $200$ proficient human teleoperated demonstrations per task. The policy receives the end-effector and object poses of the last two timesteps as observations and predicts desired changes in the end effector pose and gripper opening as actions. The operational space controller directly controls the robot in Cartesian space to follow these goals. For all other approaches, we used an inverse kinematics solver [13] to translate the desired changes in end-effector pose into desired joint state actions and control the robot in joint space. Additional implementation details can be found in Table III. For the operational space controller baseline, we used $h=8$ action execution steps and applied each action for $\Delta t=$50\text{\,}\mathrm{ms}$\$. To evaluate the impact of the safety mechanism and the inverse kinematics translation separately, we also tested the performance of our safety shield in OFF mode. Here, we compute the same intended trajectory in joint space, but never execute a failsafe trajectory.
@@
TABLE II: Constraints of our real-world tasks.
@@ VI-B Real-World Setup
Second, we conducted hardware experiments with a Franka FR3 manipulator for HRI, as safety is particularly critical in this area.
@@ VI-B1 Tasks
We considered three tasks requiring safe motion in close proximity to humans, as visualized in Figure 3:
  • Sorting: We placed two red and three green blocks randomly on the table. The robot must put the red blocks into a box while the human is taking the green blocks. We investigated different human motion patterns and speeds. To ensure safety, no collisions between the robot and the human are allowed ($c_{\text{safe},\text{SSM}}$).
@@ VI-B2 Safety Constraints
We used the SaRA framework [38] to compute reachable occupancies and performed real-time intersection checking. The safe energy thresholds in (3) have been formally identified for the different tasks [44]. Further details and numerical values are provided in Table II. We leveraged Ruckig [7] to compute the intended trajectory (5), using the standard joint limits of the FR3, velocity limits between $1.25\text{\,}$ and $2\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-1}$, an acceleration limit of $10\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-2}$ and a jerk limit of $400\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-3}$ for each joint.
@@ VI-B3 Policies
We deployed two generative policies: DP [12] using denoising diffusion [16] and SmolVLA [40] using flow matching [25]. The observations consisted of joint angles and the images of a wrist and a third-person camera, and the actions were desired changes in joint angles and the gripper opening. We provide details on the training data and the policy implementation in Table III. CRISP [33] has been used for data recording and for translating the joint commands to motor torques. The SmolVLA policy struggled with Sorting and Feeding, likely due to the high precision required; therefore, we only evaluated it on the Handover task.
@@ VI-C Real-World Results
We deployed the policies with our safety filter OFF and active, similar to the simulation experiments. As shown by the results in Table IV, the policies achieved very similar task success rates of about $80\text{\,}\%$ on average with and without our safety filter, supporting H2. The safeguarded policies do not exhibit constraint violations in our experiments. In contrast, the nominal policies are in an unsafe state $56\text{\,}\%$ of the time with each rollout containing constraint violations, resulting in a Safe Success of $0\text{\,}\%$, which confirms H1. These results highlight the necessity of online safety mechanisms when deploying DPs in dynamic environments and the effectiveness of PACS for this purpose.
TABLE IV: Impact of safeguarding in our real-world experiments.
@@
We also compare our path-consistent safety approach with a popular reactive method: control barrier functions [41]. Since control barrier functions are better suited for coexistence than for collaboration tasks, we conduct this comparison for the Sorting task. As can be seen from Table V, PACS achieved a $37\text{\,}\%$ higher success rate than the control barrier function-based filter. To further examine this inferior performance, we visualize the training data distribution and the end-effector positions of randomly selected rollouts in Figure 4. While our method kept the policy in distribution, the control barrier function often pushed the robot into out-of-distribution states from which the policy struggles to recover, leading to a higher failure rate. Based on our simulation results, we find that PACS overall achieved a $51\text{\,}\%$ higher task success rate than the safety filter using a control barrier function. Besides, we observe that policy recovery usually takes a few seconds, leading to a higher average task completion time. The average computation time of one safety step with PACS was $0.20\text{\,}\mathrm{ms}$ compared to $0.64\text{\,}\mathrm{ms}$ for the control barrier function, and recalculating the intended trajectory took about $5\text{\,}\mathrm{ms}$. This supports H4 and shows that our method can be deployed in real time in highly dynamic environments.
@@
Figure 4: End-effector paths for the Sorting task. The color intensity of the trajectories indicates the velocity, and the shaded grey areas visualize the training distribution. Our safety filter reduces the velocity without leaving the desired path when the human is nearby. In contrast, the control barrier function (CBF) pushes the robot away from unsafe states, which often leads to out-of-distribution (OOD) states from which the policy cannot recover.
@@
TABLE V: Baseline comparison and speed analysis for the Sorting task.
@@
Lastly, we tested our safety filter for the Feeding task with a real human. As shown in our accompanying video, the robot slowed down when approaching the human, and the safeguarded policy inserted and withdrew the fork carefully and precisely. These experiments highlight the potential of our approach to harness the expressiveness of DPs for highly safety-critical applications, such as healthcare.
@@ VII Conclusion
We propose PACS, a general framework for safely deploying imitation learning policies using action chunks, such as DPs and vision-language-action models, in dynamic environments. Our insight is that safety interventions should be path-consistent to avoid out-of-distribution situations and maintain high task performance. We achieve this by computing an intended trajectory from the action chunk and monitoring it using reachability analysis. Our experiments show that PACS can guarantee safety in real time and achieve high task success even on challenging real-world HRI tasks, whereas reactive safety filtering with control barrier functions frequently leads the policy into unrecoverable states. Our method is specifically designed to satisfy dynamic constraints, such as moving obstacles. Handling (semi-)static obstacles via constraint-aware online replanning is an interesting avenue for future work. We consider this work to be an important step towards deploying generative policies in safety-critical, human-centered environments.
Acknowledgements
The authors gratefully acknowledge financial support of the research training group ConVeY funded by the German Research Foundation under grant GRK 2428; financial support by the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Research, Technology and Space (BMFTR); and support by the Robotics Institute Germany, funded by BMFTR grant 16ME0997K.